OpenStudy (anonymous):

the size S of a tumor in mm cubed is given by S=2^t, where t is the number of months since the tumor was discovered a: what is the total change in the size of the tumor during the first 6 months? and b: what is the average rate of change in the size of the tumor during the first 6 months? and c: estimate the rate at which the tumor is growing at t=6?

OpenStudy (anonymous):

ick

OpenStudy (anonymous):

si!

OpenStudy (anonymous):

i got different answers than the book said so i am confused!

OpenStudy (anonymous):

at t = 0 it is 1 at t = 8 it is \[2^8=256\]

OpenStudy (anonymous):

oh it asked for t =6 it is \[2^6=64\]

OpenStudy (anonymous):

so change in size is \[64-1=63\] cubic mm

OpenStudy (anonymous):

average rate is \[\frac{63}{6}=10.5\]

OpenStudy (anonymous):

aha i forgot to subtract the one

OpenStudy (anonymous):

lol ok now right? and instantaneous rate at 6 is \[2^6\times \ln(2)\]

OpenStudy (anonymous):

how do you get the instantaneous rate? that is where i keep getting stuck...the logical progression is at am impass because i don't follow how it works.

OpenStudy (anonymous):

i got it because i know the answer. if you do not you have to approximate with things like \[\frac{2^{6.1}-2^6}{.1}\] and so on

OpenStudy (anonymous):

oh got ya! thanks

OpenStudy (anonymous):

yw